OpenStudy (anonymous):

1 multiplicative inverse _ 11

OpenStudy (dangerousjesse):

\(\large \frac{1}{11}\)?

OpenStudy (dangerousjesse):

You're basically just saying \(\frac{\large \frac{1}{1}}{11}\), which equals 11 :)

OpenStudy (anonymous):

8 - 13